# Scheduling problem of aircrafts on a same runway and dual runways

### v5

@@ V Scheduling problem of landing and takeoff aircraft on a same runway
Consider an aircraft sequence $\phi^{1}_{0}=\phi^{2}_{0}=\langle Tcf_{i},Tcf_{i+1},Tcf_{i+2}\rangle$, where the aircraft $Tcf_{i}$ and $Tcf_{i+2}$ are both takeoff (landing) aircraft, the aircraft $Tcf_{i+1}$ is a landing (takeoff) aircraft, and the separation times between the aircraft in $\phi^{1}_{0}$ and $\phi^{2}_{0}$ are different. Suppose that in $\phi_{0}^{1}$ and $\phi_{0}^{2}$, the aircraft $Tcf_{i+2}$ is relevant to either the aircraft $Tcf_{i}$ or the aircraft $Tcf_{i+1}$ and Assumption 15 holds. When $S_{i(i+1)}(\phi_{0}^{1})>S_{i(i+1)}(\phi_{0}^{2})$, then $S_{(i+1)(i+2)}(\phi_{0}^{1})=S_{(i+1)(i+2)}(\phi_{0}^{2})$ or $S_{(i+1)(i+2)}(\phi_{0}^{1})<S_{(i+1)(i+2)}(\phi_{0}^{2})$. Further, when $S_{i(i+1)}(\phi_{0}^{1})>S_{i(i+1)}(\phi_{0}^{2})$ and $S_{(i+1)(i+2)}(\phi_{0}^{1})<S_{(i+1)(i+2)}(\phi_{0}^{2})$, $S_{i(i+2)}(\phi_{0}^{2})=Y_{i(i+2)}$, i.e., in $\phi_{0}^{2}$, the aircraft $Tcf_{i+2}$ is relevant to the aircraft $Tcf_{i}$.
@@
(3) Suppose that the aircraft $Tcf_{b_{0}},Tcf_{b_{0}+2},\cdots,Tcf_{b_{0}+2m_{1}}$ are all takeoff (landing) aircraft, and the aircraft $Tcf_{b_{0}+1},Tcf_{b_{0}+3},\cdots,Tcf_{b_{0}+2m_{1}+1}$ are all landing (takeoff) aircraft, where $m_{1}\geq 0$ is an integer such that $2m_{1}+1\leq n$. If $S_{b_{0}(b_{0}+1)}(\phi_{0})\neq S_{b_{0}(b_{0}+1)}(\phi_{1})$, then $S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{0})=S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{1})$ for all $k=4,5,\cdots,m-1$.
@@ VI Scheduling problem of landing and takeoff aircraft on dual runways
(1) If the leading aircraft $Tcf_{i}$ is a takeoff aircraft and the trailing aircraft $Tcf_{j}$ is a landing aircraft, the minimum separation time $D_{P}$ is $T_{0}$, i.e., $D_{P}=T_{0}$. If the leading aircraft $Tcf_{i}$ is a landing aircraft and the trailing aircraft $Tcf_{j}$ is a takeoff aircraft, the minimum separation time $P_{D}$ is 00, i.e., $P_{D}=0$.
(2) If the two aircraft $Tcf_{i}$ and $Tcf_{j}$ are both landing or takeoff aircraft, the minimum separation time is equal to that on the same runway.
@@
Consider a group of landing aircraft and takeoff aircraft operating on dual runways. Let $\Phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n+m}\rangle$ denote the whole mixed landing and takeoff sequence on dual runway, $\Phi_{0}=\langle Tcf^{0}_{1},Tcf^{0}_{2},\cdots,Tcf^{0}_{n}\rangle$ denote the landing aircraft sequence in $\Phi$, $\Phi_{1}=\langle Tcf^{1}_{1},Tcf^{1}_{2},\cdots,Tcf^{1}_{m}\rangle$ denote the takeoff aircraft sequence in $\Phi$. From the definition of aircraft sequence, it follows that $S_{1}(\Phi)\leq S_{2}(\Phi)\leq\cdots\leq S_{n+m}(\Phi)$, $S_{1}(\Phi_{0})\leq S_{2}(\Phi_{0})\leq\cdots\leq S_{n}(\Phi_{0})$ and $S_{1}(\Phi_{1})\leq S_{2}(\Phi_{1})\leq\cdots\leq S_{m}(\Phi_{1})$.
In the following, we first discuss the aircraft relevance in Lemma 16 and Theorem 11, and based on Theorem 11, we give the optimal conditions for the optimization problem (6) in Theorem 12.
@@
Consider the aircraft sequence $\Phi$. Suppose that $Tf_{k}=[t_{0},+\infty]$ for all $k$ and Assumptions 1-4, 6-9 and 16 hold. For a landing aircraft $Tcf_{i}$ and a takeoff aircraft $Tcf_{k}$ in $\Phi$, if aircraft $Tcf_{k}$ is relevant to aircraft $Tcf_{i}$, then $k=i+1$. For a takeoff aircraft $Tcf_{i}$ and a landing aircraft $Tcf_{k}$ in $\Phi$, if aircraft $Tcf_{k}$ is relevant to aircraft $Tcf_{i}$, then $k=i+1$.
@@
Consider the aircraft sequence $\Phi$. Suppose that $Tf_{k}=[t_{0},+\infty]$ for all $k$. Under Assumptions 1-4, 6-9 and 16, the following statements hold.
@@
(3) Suppose that aircraft $Tcf_{i}$ and $Tcf_{j}$ are both landing (takeoff) aircraft, and $j-i>1$. If aircraft $Tcf_{j}$ is relevant to aircraft $Tcf_{i}$, aircraft $Tcf_{i-1}$, $Tcf_{i+1}$ and $Tcf_{j-1}$ are all takeoff (landing) aircraft.
@@
(2) From the sequence definition, $S_{h(i+1)}\geq S_{i(i+1)}\geq T_{0}$ for all $0<h<i$. If $k\neq i$, then $0<k<i$. Since $D_{P}=T_{0}$, if the landing aircraft $Tcf_{i+1}$ is relevant to a takeoff aircraft $Tcf_{k}$, then $Y_{k(i+1)}=T_{0}$. It is assumed that an aircraft lands and an aircraft takes off at the same time, it is assumed by default that the landing aircraft is ahead of the takeoff aircraft. Then the takeoff time of aircraft $Tcf_{k}$ is later than the landing time of aircraft $Tcf_{i}$, i.e., $k>i$. This yields a contradiction. Therefore, the landing aircraft $Tcf_{i+1}$ is relevant to a landing aircraft $Tcf_{k}$. From Theorem 1, $k=i$. If aircraft $Tcf_{i}$ and $Tcf_{i+1}$ are both takeoff aircraft, it can be proved that $k=i$.
(3) If one of the aircraft $Tcf_{i-1}$, $Tcf_{i+1}$ and $Tcf_{j-1}$, denoted by $Tcf_{h}$, is a landing (takeoff) aircraft, from Lemmas 1 and 7, $S_{ij}\geq Y_{ih}+Y_{hj}>Y_{ij}$, implying that aircraft $Tcf_{j}$ is not relevant to aircraft $Tcf_{i}$, which is a contradiction.
@@
(5) Since the aircraft $Tcf_{j}$ is relevant to aircraft $Tcf_{i}$, $Y_{ij}\leq 3T_{0}$. If aircraft $Tcf_{i}$ and $Tcf_{j}$ are both landing (takeoff) aircraft, $Tcf_{i+1},Tcf_{i+2},\cdots,Tcf_{j-1}$ are all takeoff (landing) aircraft. If $j-i>5$, it follows that $S_{ij}=S_{i(i+1)}+S_{(i+1)(j-1)}+S_{(j-1)j}>S_{(i+1)(j-1)}\geq 3T_{0}\geq Y_{ij}$, implying that aircraft $Tcf_{j}$ is not relevant to aircraft $Tcf_{i}$, which yields a contradiction. Suppose that one of aircraft $Tcf_{i}$ and $Tcf_{j}$ is a landing aircraft and one is a takeoff aircraft. If $j-i>5$, it is easy to see that $S_{ij}>T_{0}$, and aircraft $Tcf_{j}$ cannot be relevant to aircraft $Tcf_{i}$. Therefore, if the aircraft $Tcf_{j}$ is relevant to aircraft $Tcf_{i}$, $0<j-i\leq 5$.
@@
Theorem 11(4) shows that aircraft might be relevant to its nearest landing and takeoff aircraft ahead and is not relevant to other aircraft.
@@
Suppose that the aircraft sequence $\Phi$ is fixed. Suppose that $Tf_{k}=[t_{0},+\infty]$ for all $k$. Under Assumptions 1-4, 6-9 and 16, the following statements hold.
@@
(2) If the optimization problem (6) can be solved, the aircraft $Tcf_{2}$ is relevant to the aircraft $Tcf_{1}$ and for each $i\in\{3,4,\cdots,n+m\}$, the aircraft $Tcf_{i}$ is relevant to an aircraft $Tcf_{k}$ for $k\in\{Tcf_{i-5},\cdots,Tcf_{i-2},Tcf_{i-1}\}$.
@@
The condition of Theorem 12(2) means that aircraft $Tcf_{i}$ needs to be relevant to the nearest landing or takeoff aircraft. Moreover, it should be noted that when the subscript of a given aircraft $Tcf_{j}$ is not positive, the aircraft $Tcf_{j}$ is not taken into account by default.
Definition 10
Consider the sequence $\Phi$. If a landing/takeoff aircraft $Tcf_{i}$ is a resident-point aircraft in $\Phi_{0}/\Phi_{1}$ and relevant to a takeoff/landing aircraft in $\Phi$, the aircraft $Tcf_{i}$ is said to be a semi-resident-point aircraft in $\Phi$.
@@
Suppose that for each aircraft $Tcf_{i}\in\{2,3,\cdots,n+m\}$ in $\Phi$, the aircraft $Tcf_{i}$ is relevant to the aircraft $Tcf_{k}$ for $k\in\{Tcf_{i-5},\cdots,Tcf_{i-2},Tcf_{i-1}\}$. If aircraft $Tcf_{i}$ and $Tcf_{i-1}$ have different (takeoff/landing) operation tasks and aircraft $Tcf_{i}$ is relevant to $Tcf_{i-1}$ for $i\in\{2,3,\cdots,n+m\}$, there is at least an aircraft $Tcf_{j}\in\{Tcf_{1},Tcf_{2},\cdots,Tcf_{k-1}\}$ such that $S_{ji}(\Phi)-T_{0}<Y_{ji}$.
@@
Suppose that $Tf_{j}=[t_{0},+\infty]$ for all $j$ and Assumptions 1-4, 6-9 and 16, 17 hold.
@@
Proof: (1) In $\Phi$, $Tcf_{k+1}$ is located after $Tcf_{k}$. Thus, in $\Phi_{a}$, $Tcf_{k}$ is located after $Tcf_{k+1}$. From Remark 27 and Assumptions 16, 17, $S_{k}(\Phi_{a})=S_{k+1}(\Phi_{a})+T_{0}$. Since $Tcf_{k+1}$ is only relevant to $Tcf_{k}$ in $\Phi$, $S_{k+1}(\Phi)=S_{k}(\Phi)$. It follows that $S_{k+1}(\Phi)-S_{k+1}(\Phi_{a})+S_{k}(\Phi_{a})-S_{k}(\Phi)=T_{0}$.
@@
(1) Suppose that $S_{k}(\Phi_{a})>S_{k+1}(\Phi_{a})>S_{k-1}(\Phi_{a})$, aircraft $Tcf_{k}$ is $Tcf_{i_{0}}^{0}$ in $\Phi_{0}$ and aircraft $Tcf_{k+1}$ is $Tcf_{i_{1}}^{1}$ in $\Phi_{1}$, and $S_{i_{0}+h+1}(\Phi_{0})-S_{i_{0}+h}(\Phi_{0})=S^{a}_{i_{0}+h+1}(\Phi_{0})-S^{a}_{i_{0}+h}(\Phi_{0})$ for all $h\in\{0,1,\cdots,n-1\}$, where $S_{h}$ denotes the aircraft landing/takeoff time in $\Phi$ and $S^{a}_{h}$ denotes the aircraft landing/takeoff time in $\Phi_{a}$.
@@
(2) Suppose that $S_{k+1}(\Phi)>S_{k}(\Phi)>S_{k-1}(\Phi)$, aircraft $Tcf_{k}$ is $Tcf_{i_{1}}^{1}$ in $\Phi_{1}$ and aircraft $Tcf_{k+1}$ is $Tcf_{i_{0}}^{0}$ in $\Phi_{0}$, and $S_{i_{1}+h+1}(\Phi_{1})-S_{i_{1}+h}(\Phi_{1})=S^{a}_{i_{1}+h+1}(\Phi_{1})-S^{a}_{i_{1}+h}(\Phi_{1})$ for all $h\in\{0,1,\cdots,m-1\}$.
(2.1) If $S_{i_{0}+h}(\Phi_{0})-S_{i_{0}+h-1}(\Phi_{0})=S_{i_{1}+h+1}(\Phi_{1})-S_{i_{1}+h}(\Phi_{1})=T_{0}$ and $2T_{0}\geq S_{i_{0}+h_{0}}(\Phi_{0})-S_{i_{0}+h_{0}-1}(\Phi_{0})=S_{i_{1}+h_{0}+1}(\Phi_{1})-S_{i_{1}+h_{0}}(\Phi_{1})>T_{0}$ for all $h\in\{1,2,\cdots,h_{0}-1\}$ and some integer $0<h_{0}<\min\{n-i_{0},m-i_{1}\}$, $S^{a}_{i_{0}+h}(\Phi_{0})=S^{a}_{i_{0}+h-1}(\Phi_{0})+T_{0}$ for all $h\in\{1,2,\cdots,h_{0}-1\}$ and $S^{a}_{i_{0}+h_{0}}(\Phi_{0})=S^{a}_{i_{1}+h_{0}}(\Phi_{1})+T_{0}$.
(2.2) If $S_{i_{0}+h}(\Phi_{0})-S_{i_{0}+h-1}(\Phi_{0})=S_{i_{1}+h+1}(\Phi_{1})-S_{i_{1}+h}(\Phi_{1})=T_{0}$ and $T_{0}\leq S_{i_{0}+h_{0}}(\Phi_{0})-S_{i_{0}+h_{0}-1}(\Phi_{0})-[S_{i_{1}+h_{1}+1}(\Phi_{0})-S_{i_{1}+h_{1}}(\Phi_{0})]\leq 2T_{0}$ for all $h\in\{1,2,\cdots,h_{0}-1\}$ and some integer $0<h_{0}<\min\{n-i_{0},m-i_{1}\}$, $S^{a}_{i_{0}+h}(\Phi_{0})=S^{a}_{i_{0}+h-1}(\Phi_{0})+T_{0}$ for all $h\in\{1,2,\cdots,h_{0}-1\}$ and $S^{a}_{i_{0}+h_{0}}(\Phi_{0})=S^{a}_{i_{1}+h_{0}+1}(\Phi_{1})+T_{0}$.
@@
In Theorems 13 and 14, different orders of the aircraft $Tcf_{k}$ and $Tcf_{k+1}$ discussed might have different effects on the aircraft after them. Theorem 13 shows the relative changes of the aircraft after them and Theorem 14 shows the weakening process of the effects of the order changes of aircraft $Tcf_{k}$ and $Tcf_{k+1}$. These two theorems are also very useful in our algorithm proposed later.
Theorem 15
Proof: The proofs of Theorem 15 is obvious and omitted.
Remark 32
@@
From Theorem 16, it can be observed that the unequal separation times caused possibly by the existence of breakpoint aircraft and resident-point aircraft might result in mismatches between the landing sequence $\Phi_{0}$ and the takeoff sequence $\Phi_{1}$, but the effects of the unequal separation times would disappear after $1$ to $4$ aircraft in the absence of the influence of other breakpoint aircraft and resident-point aircraft. Moreover, it should be noted here that the forms of the minimum separation times with a sequence such as $Y_{j(j+1)}(\Phi_{1})$ and $Y_{(j+1)(j+2)}(\Phi_{1})$ are used instead of the forms $Y_{j(j+1)}$ and $Y_{(j+1)(j+2)}$ to distinguish different separation times between different sequences.

### v7

@@ V Scheduling problem of landing and takeoff aircraft on a same runway
Consider an aircraft sequence $\phi^{1}_{0}=\phi^{2}_{0}=\langle Tcf_{i},Tcf_{i+1},Tcf_{i+2}\rangle$, where the aircraft $Tcf_{i}$ and $Tcf_{i+2}$ are both takeoff (landing) aircraft, the aircraft $Tcf_{i+1}$ is a landing (takeoff) aircraft, and the separation times between the aircraft in $\phi^{1}_{0}$ and $\phi^{2}_{0}$ are different. Suppose that in $\phi_{0}^{1}$ and $\phi_{0}^{2}$, the aircraft $Tcf_{i+2}$ is relevant to either the aircraft $Tcf_{i}$ or the aircraft $Tcf_{i+1}$. When $S_{i(i+1)}(\phi_{0}^{1})>S_{i(i+1)}(\phi_{0}^{2})$, then $S_{(i+1)(i+2)}(\phi_{0}^{1})=S_{(i+1)(i+2)}(\phi_{0}^{2})$ or $S_{(i+1)(i+2)}(\phi_{0}^{1})<S_{(i+1)(i+2)}(\phi_{0}^{2})$. Further, when $S_{i(i+1)}(\phi_{0}^{1})>S_{i(i+1)}(\phi_{0}^{2})$ and $S_{(i+1)(i+2)}(\phi_{0}^{1})<S_{(i+1)(i+2)}(\phi_{0}^{2})$, $S_{i(i+2)}(\phi_{0}^{2})=Y_{i(i+2)}$, i.e., in $\phi_{0}^{2}$, the aircraft $Tcf_{i+2}$ is relevant to the aircraft $Tcf_{i}$.
@@
(3) Suppose that the aircraft $Tcf_{b_{0}},Tcf_{b_{0}+2},\cdots,Tcf_{b_{0}+2m_{1}}$ are all takeoff (landing) aircraft, and the aircraft $Tcf_{b_{0}+1},Tcf_{b_{0}+3},\cdots,Tcf_{b_{0}+2m_{1}+1}$ are all landing (takeoff) aircraft, where $m_{1}\geq 0$ is an integer such that $2m_{1}+1\leq m$. If $S_{b_{0}(b_{0}+1)}(\phi_{0})\neq S_{b_{0}(b_{0}+1)}(\phi_{1})$, then $S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{0})=S_{(b_{0}+k)(b_{0}+k+1)}(\phi_{1})$ for all $k=4,5,\cdots,m-1$.
@@ VI Scheduling problem of landing and takeoff aircraft on dual runways
(1) Suppose that $D_{P}=T_{0}$ and $P_{D}=0$, when the two aircraft have different operation tasks.
(2) Suppose that the minimum separation time is equal to that on the same runway, when the two aircraft $Tcf_{i}$ and $Tcf_{j}$ are both landing/takeoff aircraft.
@@
Consider a group of landing aircraft and takeoff aircraft operating on dual runways. Let $\Phi=\langle Tcf_{1},Tcf_{2},\cdots,$ $Tcf_{n+m}\rangle$ denote the whole mixed landing and takeoff sequence on dual runways, $\Phi_{0}=\langle Tcf^{0}_{1},Tcf^{0}_{2},\cdots,Tcf^{0}_{n}\rangle$ denote the landing aircraft sequence in $\Phi$, $\Phi_{1}=\langle Tcf^{1}_{1},Tcf^{1}_{2},\cdots,Tcf^{1}_{m}\rangle$ denote the takeoff aircraft sequence in $\Phi$. From the definition of aircraft sequence, it follows that $S_{1}(\Phi)\leq S_{2}(\Phi)\leq\cdots\leq S_{n+m}(\Phi)$, $S_{1}(\Phi_{0})\leq S_{2}(\Phi_{0})\leq\cdots\leq S_{n}(\Phi_{0})$ and $S_{1}(\Phi_{1})\leq S_{2}(\Phi_{1})\leq\cdots\leq S_{m}(\Phi_{1})$.
In the following, we first discuss the aircraft relevance in Lemma 16 and Theorem 12, and based on Theorem 12, we give an optimal conditions for the optimization problem (6) in Theorem 13.
@@
Consider the aircraft sequence $\Phi$. Suppose that $Tf_{k}=[t_{0},+\infty]$ for all $k$ and Assumptions 1-4, 6-9 and 15 hold. For a landing aircraft $Tcf_{i}$ and a takeoff aircraft $Tcf_{k}$ in $\Phi$, if aircraft $Tcf_{k}$ is relevant to aircraft $Tcf_{i}$, then $k=i+1$. For a takeoff aircraft $Tcf_{i}$ and a landing aircraft $Tcf_{k}$ in $\Phi$, if aircraft $Tcf_{k}$ is relevant to aircraft $Tcf_{i}$, then $k=i+1$.
@@
Consider the aircraft sequence $\Phi$. Suppose that $Tf_{k}=[t_{0},+\infty]$ for all $k$. Under Assumptions 1-4, 6-9 and 15, the following statements hold.
@@
(3) Suppose that aircraft $Tcf_{i}$ and $Tcf_{j}$ are both landing (takeoff) aircraft, and $j-i>1$. If aircraft $Tcf_{j}$ is relevant to aircraft $Tcf_{i}$, aircraft $Tcf_{i+1}$ and $Tcf_{j-1}$ are takeoff (landing) aircraft.
@@
(2) From the sequence definition, $S_{h(i+1)}\geq S_{i(i+1)}\geq T_{0}$ for all $0<h<i$. If $k\neq i$, then $0<k<i$. Since $D_{P}=T_{0}$, if the landing aircraft $Tcf_{i+1}$ is relevant to a takeoff aircraft $Tcf_{k}$, then $Y_{k(i+1)}=T_{0}$. It is assumed that an aircraft lands and an aircraft takes off at the same time, it is assumed by default that the landing aircraft is ahead of the takeoff aircraft. Then the takeoff time of aircraft $Tcf_{k}$ is later than the landing time of aircraft $Tcf_{i}$, i.e., $k>i$. This yields a contradiction. Therefore, the landing aircraft $Tcf_{i+1}$ is relevant to a landing aircraft $Tcf_{k}$. From Theorem 1, $k=i$. If aircraft $Tcf_{i}$ and $Tcf_{i+1}$ are both takeoff aircraft, it can be similarly proved that $k=i$.
(3) If one of the aircraft $Tcf_{i+1}$ and $Tcf_{j-1}$, denoted by $Tcf_{h}$, is a landing (takeoff) aircraft, from Lemmas 1 and 7, $S_{ij}\geq Y_{ih}+Y_{hj}>Y_{ij}$, implying that aircraft $Tcf_{j}$ is not relevant to aircraft $Tcf_{i}$, which is a contradiction.
@@
(5) Since the aircraft $Tcf_{j}$ is relevant to aircraft $Tcf_{i}$, $Y_{ij}\leq 3T_{0}$. If aircraft $Tcf_{i}$ and $Tcf_{j}$ are both landing aircraft, $Tcf_{i+1},Tcf_{i+2},\cdots,Tcf_{j-1}$ are all takeoff aircraft. If $j-i>4$, it follows that $S_{ij}\geq S_{(i+1)(j-1)}+S_{(j-1)j}\geq 3T_{0}+T_{0}>Y_{ij}$, implying that aircraft $Tcf_{j}$ is not relevant to aircraft $Tcf_{i}$, which yields a contradiction. Suppose that one of aircraft $Tcf_{i}$ and $Tcf_{j}$ is a landing aircraft and one is a takeoff aircraft. If $j-i>4$, it is easy to see that $S_{ij}>T_{0}$, and aircraft $Tcf_{j}$ cannot be relevant to aircraft $Tcf_{i}$. If aircraft $Tcf_{i}$ and $Tcf_{j}$ are both takeoff aircraft and $j-i>4$, aircraft $Tcf_{j}$ cannot be relevant to aircraft $Tcf_{i}$. Therefore, if the aircraft $Tcf_{j}$ is relevant to aircraft $Tcf_{i}$, $0<j-i\leq 4$.
@@
Theorem 12 shows that aircraft might be relevant to its nearest landing and takeoff aircraft ahead and is not relevant to other aircraft.
@@
Suppose that the orders of all aircraft in $\Phi$ is fixed. Suppose that $Tf_{k}=[t_{0},+\infty]$ for all $k$. Under Assumptions 1-4, 6-9 and 15, the following statements hold.
@@
(2) If the aircraft $Tcf_{2}$ is relevant to the aircraft $Tcf_{1}$ and for each $i\in\{3,4,\cdots,n+m\}$, and the aircraft $Tcf_{i}$ is relevant to an aircraft $Tcf_{k}$ for $k\in\{Tcf_{i-4},Tcf_{i-3},Tcf_{i-2},Tcf_{i-1}\}$, then the optimization problem (6) can be solved.
@@
Theorem 13(2) means that when each aircraft $Tcf_{i}$ is relevant to the nearest landing or takeoff aircraft, the optimization problem (6) is solved. Moreover, it should be noted that when the subscript of a given aircraft $Tcf_{j}$ is not positive, the aircraft $Tcf_{j}$ is not taken into account by default.
@@
Suppose that for each aircraft $Tcf_{i}\in\{2,3,\cdots,n+m\}$ in $\Phi$, the aircraft $Tcf_{i}$ is relevant to the aircraft $Tcf_{k}$ for some $k\in\{Tcf_{i-4},Tcf_{i-3},Tcf_{i-2},Tcf_{i-1}\}$. If aircraft $Tcf_{i}$ and $Tcf_{i-1}$ have different (takeoff/landing) operation tasks and aircraft $Tcf_{i}$ is relevant to $Tcf_{i-1}$ for $i\in\{2,3,\cdots,n+m\}$, there is at least an aircraft $Tcf_{j}\in\{Tcf_{1},Tcf_{2},\cdots,Tcf_{k-1}\}$ such that $S_{ji}(\Phi)-T_{0}<Y_{ji}$.
@@
Suppose that $Tf_{j}=[t_{0},+\infty]$ for all $j$ and Assumptions 1-4, 6-9 and 15, 16 hold.
@@
Remark 27
In the following, we will present important definitions of block subsequences, which are very useful to significantly reduce the computation to find the optimal sequence for the optimization problem (6). Before that, we need to first introduce a definition of semi-resident-point aircraft, which are key aircrafts of the block subsequences, and study the impact of the change of semi-resident-point aircraft.
Definition 10
(Semi-resident-point aircraft) Consider the sequence $\Phi$. If a landing (takeoff) aircraft $Tcf_{i}$ is a resident-point aircraft in $\Phi_{0}$ ($\Phi_{1}$) and relevant to a takeoff (landing) aircraft in $\Phi$, the aircraft $Tcf_{i}$ is said to be a semi-resident-point aircraft in $\Phi$.
Theorem 15
Consider the aircraft sequence $\Phi$. Let $\Phi_{a}$ be a sequence generated by exchanging the orders of aircraft $Tcf_{k}$ and $Tcf_{k+1}$ for some $12<k<n+m-$ 1. Suppose that Assumption 17 holds for the sequences $\Phi$ and $\Phi_{a}$, and $Tcf_{k+1}$ is relevant to only $Tcf_{k}$ in $\Phi$.
@@
Proof: (1) In $\Phi$, $Tcf_{k+1}$ is located after $Tcf_{k}$. Thus, in $\Phi_{a}$, $Tcf_{k}$ is located after $Tcf_{k+1}$. From Remark 23 and Assumptions 15, 16, $S_{k}(\Phi_{a})=S_{k+1}(\Phi_{a})+T_{0}$. Since $Tcf_{k+1}$ is relevant to only $Tcf_{k}$ in $\Phi$, $S_{k+1}(\Phi)=S_{k}(\Phi)$. It follows that $S_{k+1}(\Phi)-S_{k+1}(\Phi_{a})+S_{k}(\Phi_{a})-S_{k}(\Phi)=T_{0}$.
@@
(1) Suppose that $S_{k}(\Phi_{a})>S_{k+1}(\Phi_{a})>S_{k-1}(\Phi_{a})$, aircraft $Tcf_{k}$ is $Tcf_{i_{0}}^{0}$ in $\Phi_{0}$ and aircraft $Tcf_{k+1}$ is $Tcf_{i_{1}}^{1}$ in $\Phi_{1}$, and $S_{i_{0}+h+1}(\Phi_{0})-S_{i_{0}+h}(\Phi_{0})=S^{a}_{i_{0}+h+1}(\Phi_{0})-S^{a}_{i_{0}+h}(\Phi_{0})$ for all $h\in\{0,1,\cdots,n-i_{0}-1\}$, where $S_{h}$ denotes the aircraft landing/takeoff time in $\Phi$ and $S^{a}_{h}$ denotes the aircraft landing/takeoff time in $\Phi_{a}$.
@@
(2) Suppose that $S_{k+1}(\Phi)>S_{k}(\Phi)>S_{k-1}(\Phi)$, aircraft $Tcf_{k}$ is $Tcf_{i_{1}}^{1}$ in $\Phi_{1}$ and aircraft $Tcf_{k+1}$ is $Tcf_{i_{0}}^{0}$ in $\Phi_{0}$, and $S_{i_{1}+h+1}(\Phi_{1})-S_{i_{1}+h}(\Phi_{1})=S^{a}_{i_{1}+h+1}(\Phi_{1})-S^{a}_{i_{1}+h}(\Phi_{1})$ for all $h\in\{0,1,\cdots,m-i_{1}-1\}$.
(2.1) If $S_{i_{0}+h}(\Phi_{0})-S_{i_{0}+h-1}(\Phi_{0})=S_{i_{1}+1}(\Phi_{1})-S_{i_{1}}(\Phi_{1})=S_{i_{1}+h+1}(\Phi_{1})-S_{i_{1}+h}(\Phi_{1})=T_{0}$ and $2T_{0}\geq S_{i_{0}+h_{0}}(\Phi_{0})-S_{i_{0}+h_{0}-1}(\Phi_{0})=S_{i_{1}+h_{0}+1}(\Phi_{1})-S_{i_{1}+h_{0}}(\Phi_{1})>T_{0}$ for all $h\in\{1,2,\cdots,h_{0}-1\}$ and some integer $0<h_{0}<\min\{n-i_{0},m-i_{1}\}$, $S^{a}_{i_{0}+h}(\Phi_{0})=S^{a}_{i_{0}+h-1}(\Phi_{0})+T_{0}$ for all $h\in\{1,2,\cdots,h_{0}-1\}$ and $S^{a}_{i_{0}+h_{0}}(\Phi_{0})=S^{a}_{i_{1}+h_{0}}(\Phi_{1})+T_{0}$.
(2.2) If $S_{i_{0}+h}(\Phi_{0})-S_{i_{0}+h-1}(\Phi_{0})=S_{i_{1}+1}(\Phi_{1})-S_{i_{1}}(\Phi_{1})=S_{i_{1}+h+1}(\Phi_{1})-S_{i_{1}+h}(\Phi_{1})=T_{0}$ and $T_{0}\leq S_{i_{0}+h_{0}}(\Phi_{0})-S_{i_{0}+h_{0}-1}(\Phi_{0})-[S_{i_{1}+h_{0}+1}(\Phi_{0})-S_{i_{1}+h_{0}}(\Phi_{0})]\leq 2T_{0}$ for all $h\in\{1,2,\cdots,h_{0}-1\}$ and some integer $0<h_{0}<\min\{n-i_{0},m-i_{1}\}$, $S^{a}_{i_{0}+h}(\Phi_{0})=S^{a}_{i_{0}+h-1}(\Phi_{0})+T_{0}$ for all $h\in\{1,2,\cdots,h_{0}-1\}$ and $S^{a}_{i_{0}+h_{0}}(\Phi_{0})=S^{a}_{i_{1}+h_{0}+1}(\Phi_{1})+T_{0}$.
@@
In Theorems 15 and 16, different orders of the aircraft $Tcf_{k}$ and $Tcf_{k+1}$ discussed might have different effects on the aircraft after them. Theorem 15 shows the relative changes of the aircraft after them and Theorem 16 shows the weakening process of the effects of the order changes of aircraft $Tcf_{k}$ and $Tcf_{k+1}$. These two theorems are also very useful in our algorithm proposed later.
@@
From Theorem 17, it can be observed that the unequal separation times caused possibly by the existence of breakpoint aircraft and resident-point aircraft might result in mismatches between the landing sequence $\Phi_{0}$ and the takeoff sequence $\Phi_{1}$, but the effects of the unequal separation times would disappear after $1$ to $4$ aircraft in the absence of the influence of other breakpoint aircraft and resident-point aircraft. Moreover, it should be noted here that the forms of the minimum separation times such as $Y_{j(j+1)}(\Phi_{1})$ and $Y_{(j+1)(j+2)}(\Phi_{1})$ are used instead of the forms $Y_{j(j+1)}$ and $Y_{(j+1)(j+2)}$ to distinguish different separation times between different sequences.